# A Spatio-temporal CP decomposition analysis of New England region in the US

### v1

@@ 4.1 STPCA-CPD using Alternating Least Square method
In the following experiments we applied the STPCA-CPD algorithm to the real dataset describe in section˜2. We did the decomposition varying the number of components (i.e. the rank). We also compare the performance of our algorithm with the well known HOSVD+CPD where the initialization of the factor matrices is done using Higher Order Singular Value Decomposition (HOSVD), and a random initialization coupled with using ALS to find the CP factor matrices. We compute the reconstruction error using the relative error for each method. All the decompositions for HOSVD+CPD and random+CPD were computed using the matlab tensor toolbox (Kolda and Bader, 2009; Bader et al., 2024).
@@ 4.2 Clustering analysis of the CP factors per mode
Initialization Silhouette Scores for R =2 Best k Silhouette Scores for R = 3 Best k
HOSVD 0.4932 3 0.6528 2
Random 0.513 3 0.648 2
STPCA 0.6456 2 0.6721 2
@@ 6 Acknowledgments
All data is publicly available and cited. Code available upon request.

### v2

@@ 4.1 STPCA-CPD using Alternating Least Square method
In the following experiments we applied the STPCA-CPD algorithm to the real dataset describe in section˜2. We did the decomposition varying the number of components (i.e. the rank). We also compare the performance of our algorithm with the well known HOSVD+CPD where the initialization of the factor matrices is done using Higher Order Singular Value Decomposition (HOSVD), and a random initialization coupled with using ALS to find the CP factor matrices. We compute the reconstruction error using the relative error for each method. All the decompositions for HOSVD+CPD, STPCA+CPD, and random+CPD were computed using the matlab tensor toolbox (Kolda and Bader, 2009; Bader et al., 2024) and the tensorLab toolbox De Lathauwer et al. (2023).
@@ 4.2 Clustering analysis of the CP factors per mode
HOSVD 0.4539 3 0.6422 2
Random 0.4968 3 0.6566 2
STPCA 0.5795 2 0.6589 2
@@ 6 Acknowledgments
All data is publicly available and cited. The function stpca_to_cp_init in cp_init.R is for the initialization and is available on github.
Contact: fsanogo@bates.edu, (207) 786-6144.
Hardware requirements: No specialized hardware is required. All analyses were performed on standard personal computing hardware with sufficient RAM to handle the data size and MATLAB/R software installed.
Program language: The R and Matlab programming languages were used.
Software required: The analysis primarily utilizes the Tensor Toolbox for MATLAB (Kolda and Bader) and TensorLab (De Lathauwer et al.) within the MATLAB environment. Specific custom for the initialization was written in R. The source codes are available for downloading at the link:
https://github.com/fatsanogo/STPCA-CP-Decomposition-ALS